# Overlap-aware End-to-End Supervised Hierarchical Graph Clustering for Speaker Diarization

### v1

@@ Abstract
Speaker diarization, the task of segmenting an audio recording based on speaker identity, constitutes an important speech pre-processing step for several downstream applications. The conventional approach to diarization involves multiple steps of embedding extraction and clustering, which are often optimized in an isolated fashion. While end-to-end diarization systems attempt to learn a single model for the task, they are often cumbersome to train and require large supervised datasets. In this paper, we propose an end-to-end supervised hierarchical clustering algorithm based on graph neural networks (GNN), called End-to-end Supervised HierARchical Clustering (E-SHARC). The E-SHARC approach uses front-end mel-filterbank features as input and jointly learns an embedding extractor and the GNN clustering module, performing representation learning, metric learning, and clustering with end-to-end optimization. Further, with additional inputs from an external overlap detector, the E-SHARC approach is capable of predicting the speakers in the overlapping speech regions. The experimental evaluation on several benchmark datasets like AMI, VoxConverse and DISPLACE, illustrates that the proposed E-SHARC framework improves significantly over the state-of-art diarization systems.
@@ I Introduction
The widely used efforts in diarization attempt to segment the audio into short segments of $1-2$s, which are converted to embeddings, followed by a clustering step. The initial efforts in embedding extraction framework used Gaussian mixture modeling based background models, termed as i-vectors [1]. Recent works in speaker diarization have focused on deep learning-based models such as Time-Delay Neural Networks (TDNN) [2] and residual networks (ResNet) [3]. These models can effectively extract speaker discriminative features, called x-vectors. The embeddings are used in the clustering step with a pre-defined similarity metric, e.g., Probabilistic Linear Discriminant Analysis (PLDA) scoring [4, 5]. While this approach to diarization allows the use of pre-trained models (like embedding extraction networks and similarity scoring schemes) developed for speaker verification, the framework is highly modular, with each component optimized individually without a multi-speaker diarization loss function.
@@
The recent efforts in the development of an end-to-end neural diarization (EEND) system [10, 11] attempt to mitigate the modular issues, where the goal is to perform the entire diarization process in a single neural network. However, these models require large amounts of labeled conversational data for training and involve a large computational overhead. This can be difficult and time-consuming to obtain, and generalization to new unseen domains like multi-lingual or code-mixed settings may be challenging.
To overcome the limitations of metric learning approaches, our prior work (Singh et al. [14]) introduced a supervised hierarchical clustering approach called SHARC, which trains a graph neural network with speaker embeddings as the nodes and similarity scores as the edges. In this paper, we provide an extension of SHARC to incorporate the joint learning of the embedding extractor and the graph neural network. This approach is termed as end-to-end SHARC (E-SHARC). The framework allows the use of any deep learning-based embedding extractors, like ETDNN [27], FTDNN [28], ResNet or ECAPA-TDNN [3] models. The proposed work further tackles the problem of diarization of overlapping speech.
@@
  • Develop an end-to-end diarization system using supervised graph neural network-based clustering (E-SHARC).
  • Introduce an overlap speaker detection approach called as E-SHARC-Overlap to assign multiple speakers for the same audio region.
  • Evaluate the performance on three benchmark datasets to illustrate improvements over state-of-the-art diarization systems.
@@ II-A End-to-end neural diarization
Watanabe et al. [11] introduced a neural processing framework for speaker diarization. The goal is to train a model that predicts the speaker activity labels at each time frame in an audio recording, given the front-end time-frequency features as input. The model also attempts to identify the number of speakers in the conversation using attractor style long short-term memory (LSTM) networks [12]. Due to the difficulty in handling a large number of speakers, speaker similarity-based loss was also added to this framework [13]. In contrast, our work does not require a large amount of supervised conversational training data. Further, the E-SHARC framework can leverage large pre-trained speaker embedding models.
@@ II-B Graph clustering algorithms
The task of grouping the vertices of the graph into clusters in such a way that there are more edges within each cluster and relatively few between the clusters is called graph clustering. One of the popular methods that originated from graph theory is called spectral clustering [18]. Similarly, another algorithm called path integral clustering (PIC) [19, 8] is a graph-structural agglomerative clustering algorithm where the graph encodes the structure of the feature space. These clustering approaches are unsupervised. In contrast, we propose a supervised graph clustering framework that can be learned using the labeled data. In this work, we describe one of the first efforts in exploring hierarchical graph clustering frameworks for speaker diarization.
@@ II-C Overlap detection approaches
Medennikov et al. (2020) [20] proposed target-speaker voice activity detection (TS-VAD) to achieve accurate speaker diarization even under noisy conditions with speaker overlaps. Bullock et al. (2020) [16] proposed the Long Short-Term Memory (LSTM) based architecture for overlap detection. The model in the Pyannote [17] is trained with hand-crafted features or trainable SincNet features. The model performs 2-class classification (overlapped versus clean speech) using binary cross entropy loss. In contrast, our work utilizes overlapped speech regions in graph clustering to determine the speaker identities.
@@ III-A Background
A graph $G$ can be described by the set of vertices/nodes $V$ and edges $E$, as $G=(V,E)$. Adjacency matrix ($\bm{A}\in\mathcal{R}^{NXN}$, $N$ number of nodes) captures connections between nodes. The adjacency matrix can be obtained using similarity scores between data points. Nodes in the graph represent embeddings of individual speakers, and edges capture turn-taking between speakers in the conversation.
A graph neural network (GNN) is an effective tool for analyzing and modeling graphs. The Graph Convolution Network (GCN) [23], the most common variant of GNNs, was used in [24] for semi-supervised training using clustering output as “pseudo-labels”. Speaker diarization can be formulated as a link prediction problem [25] between speaker embeddings of segments from the same recording. The GCN is inherently transductive and does not generalize to unseen nodes. The GraphSAGE [26], another variant of GNN, is a representation learning technique suitable for dynamic graphs. Without the need for retraining, it can predict the embedding of a new node. Therefore, we have explored this variant in the current work.
III-B E-SHARC model
The proposed E-SHARC model involves two steps, namely embedding extraction and GNN-based clustering. In this work, we have explored the ETDNN model based embedding extractor.
III-C Notations
  • $\boldsymbol{\mathbb{{X}}}_{r}=\{\bm{x}_{1},...,\bm{x}_{N_{r}}\},\bm{x}_{i}\in\mathbb{R}^{D}~{}\forall i\in\{1,...,N_{r}\}$ denotes the sequence of $N_{r}$ segment embeddings for the recording $r$.
@@ III-C Notations
  • $G^{m}_{r}=(V^{m}_{r},E^{m}_{r})$ is a graph of recording $r$ at level $m$. $V^{m}_{r}$ and $E^{m}_{r}$ are the set of vertices and set of edges of the graph, respectively.
@@
  • $\boldsymbol{\mathbb{H}}^{m}_{r}=\{\bm{h}_{1}^{(m)},\bm{h}_{2}^{(m)},...,\bm{h}_{N^{m}_{r}}^{(m)}\}\in\mathcal{R}^{D^{\prime}\times N^{m}_{r}}$ denotes the node features. $N^{m}_{r}$ is the number of nodes at level $m$ for recording $r$ and $N^{0}_{r}=N_{r}$.
  • $\bm{S}^{m}_{r}\in\mathbb{R}^{N^{m}_{r}\times N^{m}_{r}}$ denotes pairwise similarity score matrix such that $[\bm{S}]_{ij}=s({\tilde{\bm{h}}_{i}}^{(m)},{\tilde{\bm{h}}_{j}}^{(m)})$, the similarity score between identity features of node $i$ and $j$.
  • $p^{m}_{ij}$ and $\hat{p}^{m}_{ij}$ denote the ground truth and predicted edge probabilities between node $i$ and $j$, respectively. at level $m$. ${P}^{m}$ and $\hat{P}^{m}$ are the ground truth and predicted edge sets, respectively, as shown in Figure 2.
  • $\hat{e}^{m}_{ij}=2\hat{p}^{m}_{ij}-1\in[-1,1]$ denotes the edge coefficient or edge weight between node $i\in\{1,..,N^{m}_{r}\}$ and $j\in J^{k}_{i}$, where $J^{k}_{i}$ represents the set of $k$-nn ($k$-nearest neighbor) nodes of $i$.
  • $\boldsymbol{\mathbb{C}}_{r}^{m}=\{\mathcal{C}_{1}^{m},...,\mathcal{C}_{n_{c}}^{m}\}$ denotes the set of clusters for level $m$ where each cluster set $\mathcal{C}_{i}^{m}$ contains all connected components/nodes without any discontinuity, $n_{c}$ is the total number of clusters formed. $N_{r}^{m+1}=n_{c}$
@@
    $k$ number of nearest-neighbors
@@
Figure 1: Block schematic of the E-SHARC algorithm with overlap handling. (a) shows E-SHARC inference containing ETDNN and GNN module for the first speaker assignment. (b) shows E-SHARC-Overlap for the second speaker assignment approach using an external overlap detector and the GNN module.
III-D Graph initialization
For both the training and inference steps of the E-SHARC framework, the first step is the creation of a graph based on the input embeddings (x-vectors). The initial graph, termed as the graph at level $m=0$, $G_{t}^{0}=(V_{t}^{0},E_{t}^{0})$, contains embeddings as the nodes in $V_{t}^{0}$ and $k$-nearest neighbor of each node based on $S_{t}^{0}$ similarity score matrix that forms the edges in $E_{t}^{0}$. The pre-trained PLDA model is used to generate similarity scores.
III-E Forward pass
Figure 1 (a) shows the block diagram of the inference step. For a test recording $t$, x-vectors $\bm{X}_{t}$ are extracted to perform hierarchical clustering. The SHARC is performed using a GNN module, which takes the graph as input and generates edge prediction probabilities.
@@ III-E Forward pass
III-E1 GNN scoring
The GNN scoring function $\Phi$ is a learnable GNN module designed for supervised clustering. The module jointly predicts node densities and edge probabilities using the input embeddings at each level. Each graph $G^{m}_{t}$, containing source and destination node pairs, is fed to the GNN scoring model. The output of the model is linkage probability $\hat{p}^{m}_{ij}$ of the nodes $v^{m}_{i}$ and $v^{m}_{j}~{}\forall i={1,..,N^{m}_{t}},j\in J^{k}_{i}$.
The edge probabilities are also used to compute node density, which measures how densely the node is connected within the cluster. A node with higher density better represents the cluster than a node with lower density. A node density using predicted edge coefficients is called pseudo density $\hat{d}^{m}_{i}$ for node $v^{m}_{i}$ and is defined as:
$\hat{d}^{m}_{i}=\frac{1}{k}\sum^{k}_{j\in J^{k}_{i}}{\hat{e}^{m}_{ij}\bm{S}^{m}_{t}(i,j)}$ (1)
The ground truth density $d^{m}_{i}$ is obtained using ground truth edge coefficient $e^{m}_{ij}=2p^{m}_{ij}-1\in\{-1,1\}$, where $p^{m}_{ij}=1$ if nodes $v^{m}_{i}$ and $v^{m}_{j}$ belong to the same cluster/speaker, otherwise $p^{m}_{ij}=0$.
III-E2 Clustering
Clustering is the process of grouping the nodes based on the presence of edge connections. After GNN scoring, clustering is performed hierarchically using the edge probabilities $\hat{p}^{m}_{ij}$ and the estimated node densities $\hat{d}_{i}^{m}$. At each level of hierarchy $m$, it creates a candidate edge set $\varepsilon(i)^{m}$, for the node $v^{m}_{i}$, with edge connection threshold $\tau$ as,
$\varepsilon(i)^{m}=\{j|(v^{m}_{i},v^{m}_{j})\in E_{t}^{m},\quad\hat{d}^{m}_{i}\leq\hat{d}^{m}_{j}\quad\textrm{and}\quad\hat{p}^{m}_{ij}\geq\tau\}$ (2)

### v2

# End-to-End Supervised Hierarchical Graph Clustering for Speaker Diarization

@@ Abstract
Speaker diarization, the task of segmenting an audio recording based on speaker identity, constitutes an important speech pre-processing step for several downstream applications. The conventional approach to diarization involves multiple steps of embedding extraction and clustering, which are often optimized in an isolated fashion. While end-to-end diarization systems attempt to learn a single model for the task, they are often cumbersome to train and require large supervised datasets. In this paper, we propose an end-to-end supervised hierarchical clustering algorithm based on graph neural networks (GNN), called End-to-end Supervised HierARchical Clustering (E-SHARC). The embedding extractor is initialized using a pre-trained x-vector model while the GNN model is trained initially using the x-vector embeddings from the pre-trained model. Finally, the E-SHARC model uses the front-end mel-filterbank features as input and jointly optimizes the embedding extractor and the GNN clustering module, performing representation learning, metric learning, and clustering with end-to-end optimization. Further, with additional inputs from an external overlap detector, the E-SHARC approach is capable of predicting the speakers in the overlapping speech regions. The experimental evaluation on benchmark datasets like AMI, Voxconverse and DISPLACE, illustrates that the proposed E-SHARC framework provides competitive diarization results using graph based clustering methods.
@@ I Introduction
Most of the diarization systems segment the audio into short segments of $1-2$s, which are converted to embeddings, followed by a clustering step. The initial efforts in embedding extraction framework used Gaussian mixture modeling based background models, termed as i-vectors [1]. Recent works in speaker diarization have focused on deep learning-based models such as Time-Delay Neural Networks (TDNN) [2, 3] and residual networks (ResNet) [4, 5]. These models can effectively extract speaker discriminative features, called x-vectors. The embeddings are used in the clustering step with a pre-defined similarity metric, e.g., Probabilistic Linear Discriminant Analysis (PLDA) scoring [6, 7]. While this approach to diarization allows the use of pre-trained models (like embedding extraction networks and similarity scoring schemes) developed for speaker verification, the framework is highly modular, with each component optimized individually without a multi-speaker diarization loss function.
@@
The recent efforts in the development of an end-to-end neural diarization (EEND) system [12, 13] attempt to mitigate the issues in the modular systems, where the goal is to perform the entire diarization process in a single neural network. However, these models require large amounts of labeled conversational data for training and involve a large computational overhead. This can be difficult and time-consuming to obtain, and generalization to new unseen domains like multi-lingual or code-mixed settings may be challenging.
To overcome the limitations of metric learning approaches, our prior work (Singh et al. [14]) introduced a supervised hierarchical clustering approach called SHARC, which trains a graph neural network with speaker embeddings as the nodes and similarity scores as the edges. End-to-end SHARC (E2E-SHARC) was proposed to incorporate the joint learning of the embedding extractor and the graph neural network. This work was inspired by Xing et al. [15], which performs supervised hierarchical clustering for images.
This paper builds on the prior work in [14] by providing thorough evaluation and analysis of the SHARC model. Further, this paper extends the E2E-SHARC model in a two-stage setting to address diarization of overlapping speech. The second stage is designed to identify the speaker identity of overlapping speaker. We refer to the overlap aware end-to-end SHARC model as E-SHARC-Overlap. The framework allows the use of any deep learning-based embedding extractors, like Extended-TDNN (ETDNN) [16], Factorized-TDNN (FTDNN) [17], ResNet [5] or ECAPA-TDNN [18] models.
@@
  • Provide detailed analysis of an end-to-end diarization system using supervised graph neural network-based clustering (E-SHARC).
  • Introduce an overlap speaker detection approach called E-SHARC-Overlap to assign multiple speakers for the same audio region.
  • Evaluate the performance on three benchmark datasets to illustrate improvements over state-of-the-art diarization systems with the oracle VAD (voice activity detection).
@@ II-A End-to-end neural diarization
Horiguchi et al. [13] introduced a neural processing framework for speaker diarization. The goal is to train a model that predicts the speaker activity labels at each time frame in an audio recording, given the front-end time-frequency features as input. The model also attempts to identify the number of speakers in the conversation using attractors based on a long short-term memory (LSTM)-based encoder-decoder network [13]. To handle large number of speakers, the modeling in [13] was applied on short chunks and the speakers across chunks were clustered using a speaker similarity matrix updated using a loss in [19]. Another approach called DiaPer [20] replaces the encoder-decoder network with a Perceiver-based decoder to incorporate cross-attention for designing the attractor in a non-autoregressive manner. It is shown to perform better than the LSTM-based attractors for CALLHOME dataset [21] in [20]. In contrast, our work does not require a large amount of supervised conversational training data. Further, the E-SHARC framework is initialized with large pre-trained speaker embedding model which is further optimized for the diarization task.
@@ II-B Graph clustering algorithms
The task of grouping the vertices of the graph into clusters in such a way that there are more edges within each cluster and relatively few between the clusters is called graph clustering. Graph clustering has been largely unsupervised like spectral clustering [22]. The widely used clustering algorithm in speaker diarization is Agglomerative Hierarchical Clustering (AHC) [23]. In AHC, the graph refers to a dendogram which illustrates the hierarchical merging process of clusters. Path integral clustering (PIC) [24, 10] is a graph-structural agglomerative clustering algorithm, where the graph encodes the structure of the feature space. In our work, we describe the first efforts in exploring hierarchical graph clustering frameworks [8, 9] for speaker diarization. We propose a supervised graph clustering framework that can be learned using the labeled data.
@@ II-C Overlap detection approaches
Recent diarization systems have relied on external overlap detection methods to perform overlap speaker assignments. Bullock et al. [25] and Sajjan et al. [26] proposed the Long Short-Term Memory (LSTM) based architecture for overlap detection. The model in Pyannote toolkit [27, 28] is trained with hand-crafted features or trainable SincNet features. The model performs speaker activity detection for $5$s chunks and then extracts homogeneous speaker segment embeddings for each chunk, which are used to cluster the segments to perform diarization. The speech activity of multiple speakers at the same time indicates overlapped speech. In contrast, our work performs graph clustering of the uniform speech segments embeddings in the entire audio during the first pass. The second pass utilizes overlapped speech regions detected by Pyannote toolkit to determine the second speaker’s identity.
@@ III Proposed Approach
In this section, we describe the proposed E-SHARC model. It involves two steps, namely embedding extraction and GNN-based clustering. In this work, we have explored the ETDNN model based embedding extractor. The GNN-based clustering (Section III-D) and GNN module training (Section III-E.2) are motivated by the work of Xing et al. [15].
@@ III-A Background
A graph $\boldsymbol{\mathbb{G}}$ can be described by the set of vertices/nodes $\boldsymbol{\mathbb{V}}$ and edges $\boldsymbol{\mathbb{E}}$, as $\boldsymbol{\mathbb{G}}=(\boldsymbol{\mathbb{V}},\boldsymbol{\mathbb{E}})$. The adjacency matrix ($\bm{A}\in\mathbb{R}^{N\times N}$, $N$ number of nodes) captures connections between nodes. The adjacency matrix can be obtained using similarity scores between data points. In the context of diarization, nodes in the graph represent embeddings of individual speakers, and edges capture the turn-taking behavior among the speakers in the conversation.
A graph neural network (GNN) is an effective tool for analyzing and modeling graphs. The Graph Convolution Network (GCN) [31], the most common variant of GNNs, was used in [32] for semi-supervised training. A GNN was used to cluster embeddings corresponding to unlabeled data and the “pseudo-labels” were used to retrain the embedding extractor. Speaker diarization can be formulated as a link prediction problem [33] between speaker embeddings of segments from the same recording. The GCN is inherently transductive and does not generalize to unseen nodes. The GraphSAGE [34], another variant of GNN, is a representation learning technique suitable for dynamic graphs. Without the need for retraining, it can predict the embedding of a new node. Therefore, we have explored this variant in the current work.
III-B Notations
Following is the list of symbols used across the paper to describe the input, output and parameters of the model. Some of the symbols will be explained in later sections based on the context.
  • $\boldsymbol{\mathbb{{X}}}_{r}=\{\bm{x}_{1},...,\bm{x}_{N_{r}}\},\bm{x}_{n}\in\mathbb{R}^{D}~{}\forall n\in\{1,...,N_{r}\}$ denotes the sequence of $N_{r}$ segment embeddings for the recording $r$.
@@ III-B Notations
  • $\boldsymbol{\mathbb{G}}^{(m)}_{r}=(\boldsymbol{\mathbb{V}}^{(m)}_{r},\boldsymbol{\mathbb{E}}^{(m)}_{r})$ is the graph of recording $r$ at level $m$. $\boldsymbol{\mathbb{V}}^{(m)}_{r}$ and $\boldsymbol{\mathbb{E}}^{(m)}_{r}$ are the set of vertices and set of edges of the graph, respectively.
@@
  • $\boldsymbol{\mathbb{H}}^{(m)}_{r}=\{\bm{h}_{1}^{(m)},\bm{h}_{2}^{(m)},...,\bm{h}_{N^{(m)}_{r}}^{(m)}\}\in\mathbb{R}^{D^{\prime}\times N^{(m)}_{r}}$ denotes the node features. Here, $D^{\prime}=2D$ is the dimension of node features, $N^{(m)}_{r}$ is the number of nodes at level $m$ for recording $r$ and $N^{0}_{r}=N_{r}$.
  • $\bm{S}^{(m)}_{r}\in\mathbb{R}^{N^{(m)}_{r}\times N^{(m)}_{r}}$ denotes pairwise similarity score matrix such that $[\bm{S}^{(m)}_{r}]_{ij}=s({\tilde{\bm{h}}_{i}}^{(m)},{\tilde{\bm{h}}_{j}}^{(m)})$, the similarity score between identity features of nodes $i$ and $j$ at level $m$.
  • $p^{(m)}_{ij}$ and $\hat{p}^{(m)}_{ij}$ denote the ground truth and predicted edge probabilities between node $i$ and $j$, respectively, at level $m$. ${P}^{(m)}$ and $\hat{P}^{(m)}$ are the ground truth and predicted edge sets, respectively, as shown in Figure 2.
  • $\hat{e}^{(m)}_{ij}=2\hat{p}^{(m)}_{ij}-1\in[-1,1]$ denotes the edge coefficient or edge weight between node $i\in\{1,..,N^{(m)}_{r}\}$ and $j\in J^{k}_{i}$, where $J^{k}_{i}$ represents the set of $k$-nn ($k$-nearest neighbor) nodes of $i$ at level $m$.
  • $\boldsymbol{\mathbb{C}}_{r}^{(m)}=\{\bm{\mathcal{C}}_{1}^{(m)},...,\bm{\mathcal{C}}_{n_{c}}^{(m)}\}$ denotes the set of clusters for level $m$ and recording $r$ where, each cluster set $\bm{\mathcal{C}}_{i}^{(m)}$ contains all connected components/nodes without any discontinuity, $n_{c}$ is the total number of clusters formed. $N_{r}^{(m+1)}=n_{c}$.
@@
    $k$ number of nearest-neighbors,
@@
    The symbols for a given recording are denoted with subscript $r$ during training or $t$ at test time. However, in some expressions below, we have not notated it explicitly.
Figure 1: Block schematic of the E-SHARC algorithm with overlap handling. (a) shows E-SHARC inference containing ETDNN and GNN modules for the first speaker assignment. (b) shows E-SHARC-Overlap for the second speaker assignment approach using an external overlap detector and the GNN module. The arrows points to the k-nearest neighbors of a node after removing the intra-cluster edges.
III-C Graph initialization
For both the training and inference steps of the E-SHARC framework, the first step is the creation of a graph based on the input embeddings (x-vectors). The initial graph, termed as the graph at level $m=0$, $\boldsymbol{\mathbb{G}}_{t}^{(0)}=(\boldsymbol{\mathbb{V}}_{t}^{(0)},\boldsymbol{\mathbb{E}}_{t}^{(0)})$, contains embeddings as the nodes in $\boldsymbol{\mathbb{V}}_{t}^{(0)}$ and $k$-nearest neighbor of each node based on $\bm{S}_{t}^{(0)}$ similarity score matrix that forms the edges in $\boldsymbol{\mathbb{E}}_{t}^{(0)}$. A pre-trained PLDA model is used to generate similarity scores.
III-D Forward pass
Figure 1 (a) shows the block diagram of the inference step. For a test recording $t$, x-vectors $\boldsymbol{\mathbb{X}}_{t}$ are extracted to perform hierarchical clustering. The SHARC is performed using a GNN module, which takes the graph as input and generates edge prediction probabilities.
@@ III-D Forward pass
III-D1 GNN scoring
The GNN scoring function $\Phi$ is a learnable GNN module designed for supervised clustering. The module jointly predicts node densities and edge probabilities using the input embeddings at each level. Each graph $\boldsymbol{\mathbb{G}}^{(m)}_{t}$, containing source and destination node pairs, is fed to the GNN scoring model. The output of the model is linkage probability $\hat{p}^{(m)}_{ij}$ of the nodes $v^{(m)}_{i}$ and $v^{(m)}_{j}~{}\forall i={1,..,N^{(m)}_{t}},j\in J^{k}_{i}$.
The edge probabilities are also used to compute node density, which measures how densely the node is connected within the cluster. A node with higher density better represents the cluster than a node with lower density. A node density using predicted edge coefficients is called pseudo density $\hat{d}^{(m)}_{i}$ for node $v^{(m)}_{i}$ and is defined as:
$\hat{d}^{(m)}_{i}=\frac{1}{k}\sum^{k}_{j\in J^{k}_{i}}{\hat{e}^{(m)}_{ij}\bm{S}^{(m)}_{t}(i,j)}$ (1)
The ground truth density $d^{(m)}_{i}$ is obtained using ground truth edge coefficient $e^{(m)}_{ij}=2p^{(m)}_{ij}-1\in\{-1,1\}$, where $p^{(m)}_{ij}=1$ if nodes $v^{(m)}_{i}$ and $v^{(m)}_{j}$ belong to the same cluster/speaker, otherwise $p^{(m)}_{ij}=0$.
III-D2 Clustering
Clustering is the process of grouping the nodes based on the presence of edge connections. After GNN scoring, clustering is performed hierarchically using the edge probabilities $\hat{p}^{(m)}_{ij}$ and the estimated node densities $\hat{d}_{i}^{(m)}$. At each level of hierarchy $m$, it creates a candidate edge set $\varepsilon(i)^{(m)}$, for the node $v^{(m)}_{i}$, with edge connection threshold $\tau$ as,
$\begin{split}\varepsilon(i)^{(m)}=&\{j|(v^{(m)}_{i},v^{(m)}_{j})\in E_{t}^{(m)},\\ &\quad\hat{d}^{(m)}_{i}\leq\hat{d}^{(m)}_{j}\quad\textrm{and}\quad\hat{p}^{(m)}_{ij}\geq\tau\}\end{split}$ (2)